Shift report 20011012 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Oct 12 2001 - 16:09:58 EDT

  • Next message: Apache: "Shift report 20011013 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Oct 12 16:09:57 2001, Shift: 08:00-16:00
    Supervisor: Michael Murray, with: none
    Log book#9 pages 143 - 145
    
    Shift Summary:
    ------------------------------------------------------------
    1) After 2 day maintanence RHIC did not come up until 15.30
    ZDC rate started at 120Hz.
    At 16.05 they called to say that they would dump the beam 
    due to a cryo problem.
    
    2) One short data run 5299, TPC pedestal run 5295, ordinary
    pedestal run 5296. Ramiero worked on DVM and the HV was 
    set to increase the gain for C1. This seemed to work.
    
    3) Several problems seemed to have disappeared because of
    work done in last 2 days. For example BRAVHO can now talk
    to C1 and H2.
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Oct 12 2001 - 16:10:35 EDT